OpenStudy (dj3strella):

Can someone help me please 1. Find the constant of variation for the quadratic variation. x 2 3 4 5 6 y 32 72 128 200 288 8 16 24 40 2. Find the constant of variation for the quadratic variation. 9y = 4x2 4/9 -2 2/3 9/4 3. If f(x) varies directly with x2, and f(x) = 75 when x = 5, find the value of f(4). 100 50 48 25 4. Write a quadratic variation equation if g(x) varies directly with x2, and g(x) = 98 when x = 7. g(x) = 7x2 g(x)= ½x2 g(x) = 14x2 g(x) = 2

OpenStudy (dan815):

quadratic has the form y=ax^2+b for the first one 32=a(2)^2+b 72=a(3)^4+b 128=a(4)^2+b.. and so on.. how can you figure out what a and b are

OpenStudy (dj3strella):

but i dont know what a and b is @dan815

OpenStudy (dan815):

you find them

OpenStudy (dan815):

you have 2 equations and 2 unknowns

OpenStudy (dan815):

or just think what number do i multiply 4 to get 32 what number do i multiply 9 to get 72. in this case b turns out to be 0, so you only have to see what a is or the constant of variation

OpenStudy (dj3strella):

8

OpenStudy (dj3strella):

@dan815